OpenStudy (roadjester):

(11.2.33) Find the area enclosed by the x-axis and the curve x=1+e^t, y=t-t^2.

OpenStudy (roadjester):

I just need the limits of integration.

OpenStudy (anonymous):

\[3x 3_{1}^{4} + \sqrt[417]{7} \cot \left[ 4 \right]\approx \sqrt{3^{2}} \approx 3\]